OpenStudy (anonymous):

I already asked this but i didn't understand the answer? Solve for a exactly: cos(2a)= -sin(a) with 0≤a≤2π

OpenStudy (anonymous):

It's multiple choice but i dont even know how you would plug them in? A. a=π B. a=π/3 C. a=π/4 D. a=π/2

OpenStudy (anonymous):

cos(2a) = 1 - 2sin^2(a) this is a standard trig identity so substituting this in the original equation 1 - 2sin^2(a) = -sin ( a) 2sin^2(a) - sin (a) - 1 = 0 this a quadratic equation which can be solved by factoring (2sin (a) + 1 )(sin (a) - 1) = 0 sin (a) = -1/2 or sin (a) = 1 angle whose sine is 1/2 = pi/6 and angle whose sine is 1 is pi/2 negative 1/2 - angle is in 3rd and 4th quadrant ie 7pi/6 and 11pi/6 there are 3 roots 7pi/6 , 11pi/6 and pi/2 in degrees this is 210,330 and 90 degrees.

OpenStudy (anonymous):

Whoops... guess jimmy got it for us. lol

OpenStudy (anonymous):

relation betewwen degrees and radians pi radians = 180 degrees 2pi radians = 360 degrees rads to degrees : - multiply by 180 / pi degrees to rads : multiply by pi / 180

OpenStudy (anonymous):

Is there anyway i can do it without trigg identities because id really like not having to try and quickly relearn them:/

OpenStudy (anonymous):

the only other way is to plug in the 4 values and see which one fits the pi/2 will satisfy the equation cos (2 * pi/2) = cos pi or cos 180 degrees; use calculator to get -1 -sin pi/2 : use calculator to find sin pi/2 or 90 degrees = 1 so -sin pi/2 = -1 so pi/2 satisfies the equation d0 is the correct answer